OpenStudy (safia21):

Write the standard form of the equation of the line described. through: (-2,4) parallel to y= -3/2x+3 through (2,4) perpendicular to y= -2/7x-5 through (5,0) perpendicular to y= -x+5

OpenStudy (katrinakaif):

Well Perpendicular lines have a negative reciprocal of their orginal slope.. and Parralllel lines have the same slope as its orginal equation.. Plug in each point after determining the slope into the equation: (y-y1) = m (x-x1)

OpenStudy (safia21):

can you show me in the equation

OpenStudy (katrinakaif):

Here I ll do one equation:

OpenStudy (katrinakaif):

First one since it says PARALLEL: the slope would be the same..so in this case M = -3/2 The ordered pair given is: (-2, 4) as in (x1, y1) Now plug them in: (y-y1)=m(x-x1) (y-4) = -3/2(x +2) y - 4 = -3/2x -3 y = -3/2x + 1 Woudl be your equation